OpenStudy (anonymous):

After an accident, police can estimate the speed that a car was traveling by measuring its skid marks. The formula r=2√5L can be used, where r is the speed, in miles per hour, and L is the length of the skid marks, in feet. Estimate the speed of a car that left skid marks (a) 50 ft long; (b) 180 ft long. How fast was the car traveling if it left skid marks 50 ft long? (Simplify your answer. Type an integer or a decimal. Round to the nearest tenth.) How fast was the car traveling if it left skid marks 160 ft long?

OpenStudy (austinl):

\(r=2\sqrt{5L}\) Is this correct?

OpenStudy (anonymous):

I believe that is the formula for the problem.

OpenStudy (austinl):

This problem is easy, you have: \((a)L=50\\ (b)L=180\) Just plug and chug! :)

OpenStudy (anonymous):

what is the problem use the formula

OpenStudy (anonymous):

take care i think L should be convered in miles

OpenStudy (austinl):

@Addu5221 it says that L is in feet right in the problem.........

OpenStudy (anonymous):

r=2√5 * 160 ?

OpenStudy (anonymous):

then u apply the formula. where do u encounter the problem?

OpenStudy (anonymous):

Not sure how to type the expression out to find the answer

OpenStudy (anonymous):

51.2 mph?

OpenStudy (austinl):

For problem a we have that the length(L) of the skid marks is 50 feet. \(r=2\sqrt{5\times50}=2\sqrt{250}=2\times15.8114=31.6228~\text{mph}\)

OpenStudy (anonymous):

31.6 = rounded to nearest tenth?

OpenStudy (austinl):

Correct. Now you try the next one!

OpenStudy (anonymous):

I am confused with your answer, is it based off of 160 ft skid ?

OpenStudy (anonymous):

That was for 50 right?

OpenStudy (austinl):

That was for the 50 foot skid. Now for the 180 foot skid :P

OpenStudy (anonymous):

its 160 not 180

OpenStudy (austinl):

Well, then plug in 160.

OpenStudy (anonymous):

I am having issues using the equation attachment

OpenStudy (austinl):

Just type it out like this 2*sqrt(###)